# The quantum Newton’s bucket: Active and passive rotations in quantum theory

### v1

@@ II.3 Rotating cylindrical wells
where $y_{Mn}$, $n=1,2,...$, is the set of solutions of the equation that satisfies the boundary conditions (see Buzzegoli and Tuchin (2023) for details), $\lambda=R\omega$, and $\kappa=\sqrt{2m\left(k_{z}^{2}/2m+|E+M\omega|\right)}$, $E+M\omega<0$, with $k_{z}$ being the wave number.
@@
II.3.1 Active rotation
For the slow rotation case, substituting Eq. (34) in equation Eq. (13),
@@ II.3.1 Active rotation
Doing the same with equation Eq. (35), the density matrix operator for the rapid rotation case is
@@
II.3.2 Passive rotation
Applying the rotation operator to both slow and rapid rotations would result in the same expression for $\hat{\rho}(t)$, since it depends only on $\hat{J}_{z}$, and not on the Hamiltonian eigenvalues. Hence, equation Eq. (19) describes the passive rotation for both situations. Thereby, clearly equations Eq. (37) and Eq. (39) are different from Eq. (19), indicating another case (a particle in a cylindrical well) in which the active and the passive rotations are not equivalent.
II.4 Coulombic potential in a 2D-rotating cylindrical well
In this section, we will evaluate one more time a Coulomb-like potential, but considering the cylindrical well previously discussed, following the same discussion in Buzzegoli and Tuchin (2023). First, the wavefunction $\psi_{0}$, which is solution to a simpler stationary problem described by
@@ II.4 Coulombic potential in a 2D-rotating cylindrical well
These energy eigenvalues are completely different from equation Eq. (15) because of the dependence of $n$ on the magnetic quantum number in the quantization condition, which drastically changes the discussion of the equivalence between active and passive rotations in the presence of a potential $1/r$.
II.4.1 Active rotation
Using Eq. (41) in the equation Eq. (13), we have
@@ II.4.2 Passive rotation
III General principles: symmetries and equilibrium
The problems above are important as it illustrates an important difference between classical and quantum dynamics. In the first case, cylindrical symmetry and spherical symmetry are equivalent for rotations around the axis of the cylinder (we’ll call it $z$ here), because there is no fundamental uncertainty related to the commutators between generators.
In the quantum case, however, wavefunctions are always smeared in accordance to the uncertainty principle. For spherical-type symmetries, this symmetry is bound by the Wigner-Eckart theorem Sakurai and Napolitano (2020), which makes the density matrix factorize into a part depending only on the Casimir and the rest.
@@ III General principles: symmetries and equilibrium
For spatial symmetries, a representation of this is spherical harmonics, while for spin one can use SU(2) matrices ($SO(3)\equiv SU(2)/Z_{2}$). Cylindrical symmetry is described by the group $SO(2)\times T_{1}$, where $T_{1}$ are translations in the $z$ direction. This symmetry group shares some, but not all generators with the spherical group, and consequently, the Casimir is different ($J_{x}^{2}+J_{y}^{2}$). In spacetime, its representation is given by Bessel functions.
@@
For a passive rotation, Eq. (43) will always be true as long as the system is rotationally invariant for that axis. But for an active one, it will depend on the symmetries of the potential causing the rotation, and if this potential breaks just one of the generators of the rotation group (e.g. if the potential has cylindrical rather than spherical symmetry), the factorization of Eq. (43) will break down. In this case, passive and active rotations will obviously not be equivalent, since, in one case, there is favorization and, in the the other, there is not.
@@
Thus, the equivalence of active and passive rotations requires that all generators of the rotation group to commute, even if those are irrelevant to the physical rotation. This is an inherently quantum effect, related intimately to the uncertainty due to non-commuting operators included in the wavefunctions.
@@
This reasoning has a parallel in the famous result of N. Bohr’s doctoral research van Vleck (1934) known as the Bohr–Van Leeuwen theorem. The vanishing of classical magnetization can be thought of as a consequence of the fact that the energy of the system is degenerate w.r.t. the magnetic field (which only acts on the angular momentum), and by the symmetries of spacetime, the partition function only depends on energy. In other words, since the system is invariant under rotations of the axis parallel to the magnetic field, but the only effect of the magnetic field is to induce time-dependent rotation along that axis of each particle, active and passive rotations coincide and since classically passive rotations are irrelevant, so is the magnetic field. In the quantum regime, the presence of the magnetic potential in the momentum implies $[p_{x},y]\neq 0$. Even though, the system is only rotating in the $z$-axis and the wave-function structure is “stretched” in the $x,y$ axis as well (the stretching direction is Gauge-dependent, but its extent is not). This changes $p^{2}$ and hence $\left\langle H\right\rangle$.
@@
What happens if, in addition, the system is in global equilibrium? In such case, the velocity of each energy cell can only be a killing vector Becattini et al. (2021b); Palermo et al. (2021). As examined in Selch et al. (2023), the equivalence of active and passive rotations trivially follows provided the underlying Hamiltonian is Lorentz invariant. The reason is that the maximally mixed state will become another maximally mixed state when translated by a Killing vector. Thus, the second part of Eq. (4) is irrelevant and the first term will be satisfied by default by the values of temperature and polarizability, which satisfy the maximum entropy condition of Becattini et al. (2021b); Palermo et al. (2021). Instead of Eq. (43), we have the killing vector condition and maximal mixedness in both passive and active rotations.
@@
Perfect local equilibrium Dubovsky et al. (2012), which implies ergodicity of the microscopic degrees of freedom in every infinitesimal cell Torrieri (2024a), will give an analogous result provided the ergodic hypothesis is satisfied exactly for every cell, since the action on a maximally mixed state preserves maximal mixedness. Away from these highly unrealistic limits, though, a violation of Eq. (4) will generally mean that active and passive rotations should be different.
@@
In the Newtonian limit, the problem reduces to the one in section  (II.1), where, as we saw, the passive and active rotations coincide. For the relativistic linearized limit, as shown in Feynman (1996), the tensorial structure of the theory, together with the Gauge consistency condition forces the interaction lagrangian to be of the form
@@
where $T_{\mu\nu}\sim\frac{1}{\sqrt{\mathrm{det}\left(\eta_{\mu\nu}+h_{\mu\nu}\right)}}\frac{\delta L}{\delta h^{\mu\nu}}$ is the energy momentum tensor, of the field, given by the right represtentation of the SU(2) group and $h_{\mu\nu}$ the linear metric perturbation. As shown in Feynman (1996) it follows that the motion of a particle is a geodesic, with time-dilation given by factors of $\epsilon h_{\mu\nu}dx^{\mu}dx^{\nu}$ where $dx_{\mu}$ is a worldline element,
[…]
For rotating motion (actually for general motion in gravitational fields), this automatically satisfies Eq. (4), since the generators of the rotation also generate the metric giving rise to the rotation and time dilation appears in the “active” rotation (with $h_{\mu\nu}$ and the passive one (in the freely falling frame) in the same way. Thus, for general relativity active and passive motion in a gravitational field are locally equivalent, which is in fact one of the ways to see the equivalence principle.
IV Field theory
IV.1 The quantum fluctuation in an electromagnetic field of electrons in rings
The Hamiltonian that takes into consideration the couple between spin and angular rotation due the electron rotation in a ring is given by:
$\hat{H}=\frac{1}{2}\hbar\bm{\Omega}\cdot\bm{\hat{\sigma}}$ (46)
where we can write the angular rotation in the ring in two components: the first classical $\bm{\omega}$ and the second referent to quantum fluctuation $\delta\bm{\omega}$. Thus Bell and Leinaas (1987),
@@ IV.1 The quantum fluctuation in an electromagnetic field of electrons in rings
where the quantum fluctuation is defined as Bell and Leinaas (1987):
$\delta\omega_{\pm}=-\frac{e}{2mc}\left(gB^{\prime}_{f\pm}+2E^{\prime}_{fz}\right)$ (48)
$B^{\prime}_{f\pm}$ and $E^{\prime}_{fz}$ are magnetic and electric free fields. One possible interpretation is that these fields can be understood as the fields created due to the acceleration of the electron in the ring.
The quantum fluctuation, $\delta\bm{\omega}$, occurs only in the active rotation, and there will be no influence of quantum fluctuation on the passive rotation. Thus, we will see that active and passive rotations differ in these scenarios. The possible observable is the electron polarization effect in both active and passive rotations.
To understand this system it helps remembering the famous “Feynman’s disk paradox”(Feynman et al., 1963, Ch. 17-4): A disk with static bead-like charges near its perimeter can freely rotate around its axis. The disk has a strong magnetic field parallel to its rotation axis, generated by a superconducting current. Initially, the system is at rest. The disk is then allowed to heat up until the magnetic field collapses, which generates a strong circular electric field tangent to the disk’s perimeter due to the presence of the charges. This would result in a net torque and, consequently, in the rotation of the disk. However, a paradox arises when considering that the angular momentum of the system must be conserved. Therefore, there should be no rotation, since the disk is initially at rest.
The resolution of this paradox has to do with the angular momentum contained in the electromagnetic field, which of course will, at infinity, contain radiated electromagnetic waves. Such asymptotic states will exist for active rotations but not for passive ones (where they will appear as zero-momentum modes, as seen in Higuchi et al. (1992)). In the setup of Bell and Leinaas (1987) spin flips will be dynamically carried out by the emission and absorption of photons, which only happens when active rotations are performed.
IV.2 From quantum mechanics to quantum field theory
We can now say something about quantum fields, relevant to the problems such as Braguta et al. (2023); Chernodub (2021); Ayala et al. (2021); Prokhorov et al. (2020); Yamamoto and Hirono (2013); Aristova et al. (2016); Wang et al. (2019) and calculations such as Kim et al. (2023). Of course, the original problem of Newton is also a (classical) field theory problem, since water in a bucket is a classical field. Let us therefore say something about this problem in the language developed throghout this work. The Hamiltonian for the system is
@@ IV.2 From quantum mechanics to quantum field theory
where the first term is a simple rigid body rotating hamiltonian, , but the subsequent terms contain a “large” (call it $N$) number of degrees of freedom
@@
For $N\rightarrow\infty$ and no dissipation $V_{water}$ converges to the lagrangian of incompressible fluid dynamics, famously represented via volume preserving diffeomorphisms Arnold (1966). By inspection the potential term $V_{contact}$ and $V_{grav}$ and the generators of the diffeomorphism group of Arnold (1966) do not commute with $J_{z}$ so we do not expect active and passive rotations to be the same even in the classical limit, thereby confirming Newton (1687).
Let us now see what happens in quantum field theory. The limit of QM from QFT is Zee (2003)
@@
are given by the expectation value of Wilson line operators between sources $W(J_{1},J_{2},...)$
Deriving potentials for generic fields, especially in the strongly coupled case, is of course a complicated task, but the sum rules in Kim et al. (2023) give us some clue. As argued in that work, one should follow the decomposition of Ji (1997) into a quark spin $S_{q}$ and a quark and gluon angular momenta $L_{q},L_{k}$. The first is local, the other two are not. If a decomposition can be found of $W(J_{1},J_{2},...)$ into the relevant $S_{q}$,$L_{q,k}$ the latter will generally lead to differences in active and passive rotation.
@@
The spin term is local and the other terms are not. Thus the only the first term should be invariant between passive and active rotations. Constructing a Polyakov loop from $W(...)$ and making such an operator product expansion from the Polyakov loop should be sensitive to the difference between active and passive rotations of thermal media.
@@
Of course, the identification of QFT correlators with quantum potentials only works for time-invariant problems. Thus, quantum field theory can have “imaginary” values for the Wilson line operators, which signals an instability of the state in question Peskin and Schroeder (1995).
@@
Since intuitively only active rotations can cause a state to break De Moura et al. (2023b); Liang and Lin (2025); Chen et al. (2021), and bound states break via tidal forces, perhaps the imaginary part of $W(...)$ provides a QFT “simple” criterion that active and passive rotations are not in fact the same.
Such considerations will be examined in future works. As a complication, condensates do seem to depend on passive rotations Braguta et al. (2023) so perhaps, because of operator non-commutativity, observing the system via rotation does change it’s microscopic structure. Such questions can perhaps be clarified by considering a Gibbsian coarse-grained entropy rather than a Boltzmannian one for equilibrium (Dore et al. (2022) and ongoing work on Gibbsian hydrodynamics).
V Conclusions

### v2

@@ 2.3 Rotating cylindrical wells
where $y_{Mn}$, $n=1,2,$..., is the set of solutions of the equation that satisfies the boundary conditions (see [Buzzegoli and Tuchin(2023)] for details), $\lambda=R\omega$, and $\kappa=\sqrt{2m\left(k_{z}^{2}/2m+|E+M\omega|\right)}$, $E+M\omega<0$, with $k_{z}$ being the wave number.
@@
2.3.1 Active rotation
For the slow rotation case, substituting Eq. (34) in Eq. (13),
@@ 2.3.1 Active rotation
Doing the same with Eq. (35), the density matrix operator for the rapid rotation case is
@@
2.3.2 Passive rotation
Applying the rotation operator to both slow and rapid rotations would result in the same expression for $\hat{\rho}(t)$, since it depends only on $\hat{J}_{z}$, and not on the Hamiltonian eigenvalues. Hence, Eq. (19) describes the passive rotation for both situations. Thereby, clearly Eq. (37) and Eq. (39) are different from Eq. (19), indicating another case (a particle in a cylindrical well) in which the active and the passive rotations are not equivalent.
2.4 Coulombic potential in a 2D-rotating cylindrical well
In this section, we will evaluate one more time a Coulomb-like potential, but considering the cylindrical well previously discussed, following the same discussion in [Buzzegoli and Tuchin(2023)]. First, the wavefunction $\psi_{0}$, which is solution to a simpler stationary problem described by
@@ 2.4 Coulombic potential in a 2D-rotating cylindrical well
These energy eigenvalues are completely different from Eq. (15) because of the dependence of $n$ on the magnetic quantum number in the quantization condition, which drastically changes the discussion of the equivalence between active and passive rotations in the presence of a potential proportional to $1/r$.
2.4.1 Active rotation
Using Eq. (41) in Eq. (13), we have
@@ 2.4.2 Passive rotation
3 General principles: symmetries and equilibrium
The problems above are important as they illustrate an important difference between classical and quantum dynamics. In the first case, cylindrical symmetry and spherical symmetry are equivalent for rotations around the axis of the cylinder (we’ll call it $z$ here), because there is no fundamental uncertainty related to the commutators between generators.
In the quantum case, however, wavefunctions are always smeared in accordance to the uncertainty principle. For spherical-type symmetries, this symmetry is bounded by the Wigner-Eckart theorem [Sakurai and Napolitano(2020)], which makes the density matrix factorize into a part depending only on the Casimir and the rest.
@@ 3 General principles: symmetries and equilibrium
For spatial symmetries, a representation of this is spherical harmonics, while for spin, one can use SU(2) matrices ($SO(3)\equiv SU(2)/Z_{2}$). Cylindrical symmetry is described by the group $SO(2)\times T_{1}$, where $T_{1}$ are translations in the $z$ direction. This symmetry group shares some, but not all generators with the spherical group, and consequently, the Casimir is different ($J_{x}^{2}+J_{y}^{2}$). In spacetime, its representation is given by Bessel functions.
@@
For a passive rotation, Eq. (43) will always be true as long as the system is rotationally invariant for that axis. But for an active one, it will depend on the symmetries of the potential causing the rotation, and if this potential breaks just one of the generators of the rotation group (e.g. if the potential has cylindrical rather than spherical symmetry), the factorization of Eq. (43) will break down. In this case, passive and active rotations will obviously not be equivalent, since, in one case, there is factorization and, in the the other, there is not.
@@
Thus, the equivalence of active and passive rotations requires that all generators of the rotation group to commute, even if those are irrelevant to the physical rotation. This is an inherently quantum effect, related intimately to the uncertainty (expressed via cumulants of operators) due to non-commuting operators included in the wavefunctions: if a generator does not commute with the Hamiltonian, even one perpendicular to the rotation axis, the active rotation will introduce an uncertainty in the value of that operator, which will be absent in the passive rotation. Since such uncertainties are measurable, active and passive rotations will not be the same.
@@
This reasoning has a parallel in the famous result of N. Bohr’s doctoral research [van Vleck(1934)] known as the Bohr–Van Leeuwen theorem. The vanishing of classical magnetization can be thought of as a consequence of the fact that the energy of the system is degenerate w.r.t. the magnetic field (which only acts on the angular momentum), and by the symmetries of spacetime, the partition function only depends on energy. In other words, since the system is invariant under rotations of the axis parallel to the magnetic field, but the only effect of the magnetic field is to induce time-dependent rotation along that axis of each particle, active and passive rotations coincide and, since classically passive rotations are irrelevant, so is the magnetic field. In the quantum regime, the presence of the magnetic potential in the momentum implies $p_{x},y]\neq 0$. Even though, the system is only rotating in the $z$-axis and the wave-function structure is “stretched” in the $x,y$ axis as well (the stretching direction is Gauge-dependent, but its extent is not). This changes $p^{2}$ and, hence, $\left\langle H\right\rangle$.
@@
What happens if, in addition, the system is in global equilibrium? In such case, the velocity of each energy cell can only be a killing vector [Becattini et al.(2021b)Becattini, Buzzegoli, and Palermo, Palermo et al.(2021)Palermo, Buzzegoli, and Becattini]. As examined in [Selch et al.(2023)Selch, Abramchuk, and Zubkov], the equivalence of active and passive rotations trivially follows provided the underlying Hamiltonian is Lorentz invariant. The reason is that the maximally mixed state will become another maximally mixed state when translated by a Killing vector. Thus, the second part of Eq. (4) is irrelevant and the first term will be satisfied by default by the values of temperature and polarizability, which satisfy the maximum entropy condition of [Becattini et al.(2021b)Becattini, Buzzegoli, and Palermo, Palermo et al.(2021)Palermo, Buzzegoli, and Becattini]. Instead of Eq. (43), we have the killing vector condition and maximal mixedness in both passive and active rotations (in fact, the rigid body approximation is nothing else but the Killing vector criterion in the Galilean symmetry limit).
@@
Perfect local equilibrium [Dubovsky et al.(2012)Dubovsky, Hui, Nicolis, and Son], which implies ergodicity of the microscopic degrees of freedom in every infinitesimal cell [Torrieri(2024a)], will give an analogous result provided the ergodic hypothesis is satisfied exactly for every cell, since the action on a maximally mixed state preserves maximal mixedness. Away from these highly unrealistic limits, though, a violation of Eq. (4) will generally mean that active and passive rotations should be different.
@@
In the Newtonian limit, the problem reduces to the one in section  (2.1), where, as we saw, the passive and active rotations coincide. For the relativistic linearized limit, as shown in [Feynman(1996)], the tensorial structure of the theory, together with the Gauge consistency condition forces the interaction Lagrangian to be of the form
@@
where $T_{\mu\nu}\sim\frac{1}{\sqrt{\mathrm{det}\left(\eta_{\mu\nu}+h_{\mu\nu}\right)}}\frac{\delta L}{\delta h^{\mu\nu}}$ is the energy momentum tensor of the field, given by the right representation of the SU(2) group, and $h_{\mu\nu}$ the linear metric perturbation. As shown in [Feynman(1996)], it follows that the motion of a particle is a geodesic, with time-dilation given by factors of $\epsilon h_{\mu\nu}dx^{\mu}dx^{\nu}$, where $dx_{\mu}$ is a worldline element. Hence,
$S=\int d^{4}xL_{gr}\simeq\int(1+\epsilon)dtd^{3}xL_{matter}+\mathcal{O}\left(h^{2},\partial^{2}h\right).$ (45)
For rotating motion (actually, for general motion in gravitational fields), this automatically satisfies Eq. (4), since the generators of the rotation also generate the metric, giving rise to the rotation and time dilation appears in the “active” rotation (with $h_{\mu\nu}$ and the passive one (in the freely falling frame) in the same way). Thus, for general relativity, active and passive motion in a gravitational field are locally equivalent, which is, in fact, one of the ways to see the equivalence principle.
4 Field theory
The discussion in the previous section assumed “static” electric and magnetic fields in every frame, which is, at best, an approximation in the passive frame (the magnetic field in Eq. (20) will create a time-varying electric field, and indeed the ambiguity of which field is which to which observer was a motivator for relativity [Einstein(1905)]). Deviating from this approximation will however inevitably result, at the same order in time gradients, in the creation of electromagnetic waves, forcing us to consider a field theory rather than a mechanical system if interactions are present [Currie et al.(1963)Currie, Jordan, and Sudarshan].
When one goes from quantum mechanics to quantum field theory, the infinity of degrees of freedom and the presence of inequivalent representations lead to the usual mathematical ambiguities. Nevertheless, the definitions Eq. (1) and Eq. (3) can be adapted straight-forwardly to a quantum field theory using the path integral method [Peskin and Schroeder(1995)]. Since the active rotation is defined in the semi-classical limit, the background field method can be used to provide a suitable definition. For electromagnetism, this means
$\hat{p}_{0}\rightarrow\hat{p}_{0}-eA_{0}\phantom{AA},\phantom{AA}A_{0}=V\hat{x}$ (46)
which can be straight-forwardly solved as it does not alter the Gaussian nature of the integral.
The passive rotation has to be incorporated by a source term in the detector,
$S=\int d^{4}xL(A,\psi)\rightarrow\int d^{4}x\left[L(A,\psi)+J_{\mu}(x,t)A^{\mu}\right].$ (47)
Systematically working out the consequences of these is involved and has to be done on a case by case basis, but it is clear they continue, in general, to be different. In the next few sections, we shall examine a few simple cases and try to generalize the implications.
4.1 The quantum fluctuation in an electromagnetic field of electrons in rings
We are now ready to go beyond the test field assumptions of the previous section, and examine what happens when the passively rotating observer sees the static magnetic field of Eq. (20).
The Hamiltonian that takes into consideration the coupling between spin and angular rotation due the electron rotation in a ring is given by
$\hat{H}=\frac{1}{2}\hbar\bm{\Omega}\cdot\bm{\hat{\sigma}}$, (48)
where we can write the angular rotation in the ring in two components: the first classical $\bm{\omega}$ and the second referent to quantum fluctuation $\delta\bm{\omega}$. Thus [Bell and Leinaas(1987)],
@@ 4.1 The quantum fluctuation in an electromagnetic field of electrons in rings
where the quantum fluctuation is defined as [Bell and Leinaas(1987)]:
$\delta\omega_{\pm}=-\frac{e}{2mc}\left(gB^{\prime}_{f\pm}+2E^{\prime}_{fz}\right).$ (50)
$B^{\prime}_{f\pm}$ and $E^{\prime}_{fz}$ are magnetic and electric free fields, respectively. One possible interpretation is that these fields can be understood as the fields created due to the acceleration of the electron in the ring.
The quantum fluctuation, $\delta\bm{\omega}$, occurs only in the active rotation, and there will be no influence of quantum fluctuation on the passive rotation. Thus, we will see that active and passive rotations differ in these scenarios. The possible observable is the electron polarization in both active and passive rotations.
To understand this system, it is helpful to remember the famous “Feynman’s disk paradox”[Feynman et al.(1963)Feynman, Leighton, and Sands, Ch. 17-4]: a disk with static bead-like charges near its perimeter can freely rotate around its axis. The disk has a strong magnetic field parallel to its rotation axis, generated by a superconducting current. Initially, the system is at rest. The disk is then allowed to heat up until the magnetic field collapses, which generates a strong circular electric field tangent to the disk’s perimeter due to the presence of the charges. This would result in a net torque and, consequently, in the rotation of the disk. However, a paradox arises when considering that the angular momentum of the system must be conserved. Therefore, there should be no rotation, since the disk is initially at rest.
The resolution of this paradox has to do with the angular momentum contained in the electromagnetic field, which of course, will, at infinity, contain radiated electromagnetic waves (a manifestation of the theorem [Currie et al.(1963)Currie, Jordan, and Sudarshan]). Such asymptotic states will exist for active rotations, but not for passive ones (where they will appear as zero-momentum modes, as seen in [Higuchi et al.(1992)Higuchi, Matsas, and Sudarsky]). In the setup of [Bell and Leinaas(1987)], spin flips will be dynamically carried out by the emission and absorption of photons, which only happens when active rotations are performed.
4.2 Revisiting the bucket and general field theory lessons
We can now say something about quantum fields, relevant to the problems such as [Braguta et al.(2023)Braguta, Kudrov, Roenko, Sychev, and Chernodub, Chernodub(2021), Ayala et al.(2021)Ayala, Hernández, Raya, and Zamora, Prokhorov et al.(2020)Prokhorov, Teryaev, and Zakharov, Yamamoto and Hirono(2013), Aristova et al.(2016)Aristova, Frenklakh, Gorsky, and Kharzeev, Wang et al.(2019)Wang, Wei, Li, and Huang] and calculations such as [Kim et al.(2023)Kim, Cho, and Lee]. Of course, the original problem of Newton is also a (classical) field theory problem, since water in a bucket is a classical field. Let us, therefore, say something about this problem in the language developed throughout this work. The Hamiltonian for the system is
@@ 4.2 Revisiting the bucket and general field theory lessons
where the first term is a simple rigid body rotating Hamiltonian, but the subsequent terms contain a “large” (call it $N$) number of degrees of freedom,
@@
For $N\rightarrow\infty$ and no dissipation, $V_{water}$ converges to the Lagrangian of incompressible fluid dynamics, famously represented via volume preserving diffeomorphisms [Arnold(1966)]. By inspection, the potential terms $V_{contact}$ and $V_{grav}$ and the generators of the diffeomorphism group of [Arnold(1966)] do not commute with $J_{z}$ (and more in general, the Galileo group), so we do not expect active and passive rotations to be the same even in the classical limit, thereby confirming [Newton(1687)].
Let us now see what happens in quantum field theory. The limit of QM from QFT is [Zee(2003)]
@@
[…]
@@
The spin term is local and the other terms are not. Thus, the only the first term should be invariant between passive and active rotations. Constructing a Polyakov loop from W()W(...) and making such an operator product expansion from the Polyakov loop should be sensitive to the difference between active and passive rotations of thermal media.
@@
Of course, the identification of QFT correlators with quantum potentials only works for time-invariant problems. Thus, quantum field theory can have “imaginary” values for the Wilson line operators, which signals an instability of the state in question [Peskin and Schroeder(1995)].
@@
Since intuitively only active rotations can cause a state to break [De Moura et al.(2023b)De Moura, Goncalves, and Torrieri, Liang and Lin(2025), Chen et al.(2021)Chen, Zhao, and Zhuang], and bound states break via tidal forces, perhaps the imaginary part of W()W(...) provides a QFT “simple” criterion that active and passive rotations are not in fact the same.
An additional complication, with likely relevance to the ambiguities pointed out in [Ohsaku(2004), Benic and Fukushima(2015), Salluce et al.(2024)Salluce, Pasini, Flachi, Pittelli, and Ansoldi, Casado-Turrión and Dobado(2019), Chernodub(2025), Unruh and Weiss(1984)], is that quantum field theory, because of the infinite number of degrees of freedom, does not meet the axioms of the Stone Von Neumann theorem, on which the symmetry arguments in section  (3) are based. The fact that the vacuum [De Paola and Svaiter(2000)] does not, in general, have the symmetries of the Lagrangian is a direct consequence of this, and it can occur in several different ways, such as spontaneously broken symmetries (the breaking of the symmetries of the Lagrangian by infrared modes), anomalously broken symmetries (an interplay between infrared and ultraviolet modes via counter-terms) and boundary conditions (the so called Unruh and Hawking effects). In setups such as [Ohsaku(2004), Benic and Fukushima(2015), Salluce et al.(2024)Salluce, Pasini, Flachi, Pittelli, and Ansoldi, Casado-Turrión and Dobado(2019), Chernodub(2025), Unruh and Weiss(1984)], at least two of the effects numerated above are present, leading to ambiguity.
This issue will be examined quantitatively in a future work. We note that the equivalence of active and passive transformations inevitably implies that the effect of accelleration is indistinguishable from that of thermal baths, which some of [Ohsaku(2004), Benic and Fukushima(2015), Salluce et al.(2024)Salluce, Pasini, Flachi, Pittelli, and Ansoldi, Casado-Turrión and Dobado(2019), Chernodub(2025), Unruh and Weiss(1984)] disagree with. On the other hand, because of local operator non-commutativity, observing the system via an accellerated or rotating detector could in principle change it’s microscopic structure. We note however that the systems examined in [Ohsaku(2004), Benic and Fukushima(2015), Salluce et al.(2024)Salluce, Pasini, Flachi, Pittelli, and Ansoldi, Casado-Turrión and Dobado(2019), Chernodub(2025), Unruh and Weiss(1984)] have in a sense two temperatures, the internal one of the accellerating system and accelleration. In such a case, given the arguments of section  (3) that equilibrium is what determines the relation between passive and active transformations, one must think carefully as to what is meant by equilibrium. Considering a Gibbsian coarse-grained entropy rather than a Boltzmannian one for equilibrium ([Dore et al.(2022)Dore, Gavassino, Montenegro, Shokri, and Torrieri] and ongoing work on Gibbsian hydrodynamics) might clarify whether active and passive non-inertial transformations coincide for a quantum field with multiple vacua.
5 Conclusions